OpenStudy (anonymous):

The first three terms of a sequence are given. 11,16,21,... Find the 42nd term.

OpenStudy (cwrw238):

this looks like an arithmetic series where common difference = 5 ( 16-11 = 5 and 21-16 = 5 nth term = a + d(n-1) where a = first term and d = common difference so 42d term = 11 + 5(42-1)